# An Entropy-based Text Watermarking Detection Method

### v2

@@ 6.2 Main Results
Table 1 shows that for the low-entropy scenario, code detection, our detector would achieve the best detection accuracy against other baselines. Specifically, for HumanEval and MBPP respectively, our method could achieve a 10.6% and 5.6% improvement than SWEET in true positive rate while maintaining a low-than 5% false positive rate.
Table 1 shows that for the high-entropy scenarios using Rotowire and C4 datasets, our method is better than SWEET and is very similar to the original KGW method, with very minor differences.
@@ 6.3 Analysis
Impact of Entropy-based Weights. In figure 2, we gave a few visualizations to show the reason why EWD surpassed other baselines in detection performance. Figure 2(a) shows the z-scores of both watermarked and human texts in different datasets, detected by different methods. Our EWD method would result in overall higher z-score for watermarked texts, and slightly lower z-score for human texts. The z-score increase is more significant in low-entropy code generation tasks than in the Rotowire dataset. By enlarging the gap between the z-scores of watermarked and human texts, our detector could better distinguish the watermarked texts. In figure 2(b), we plotted the relationship of a token’s entropy-based weight and probability of that token being watermarked. In watermarked text, the probability is proportional to the token weight. That is to say, the tokens assigned with a large weight are very likely to be watermarked and contribute to a larger z-score. On the contrary, in human-written texts, there doesn’t exist such a relationship, explaining why the z-scores of human text remain low.
Ablation Study. In the experiments, we utilized a linear function to gain the entropy-based weights for each token. We also studied a variety of monotonic increasing functions as shown in figure 3. Specifically, each function’s gradient is different. For our custom Sigmoid functions, the gradient starts off high and gradually decreases, that is to say, tokens with relatively low entropy will have significant differences in weights, while tokens with relatively high entropy will exhibit smaller differences. Similarly for our custom Exponential functions, the gradient starts off low and gradually increases, meaning that tokens with relatively low entropy will have smaller differences in weight, while tokens with relatively high entropy will exhibit larger differences. We use function strength to represent how far the function deviates from being linear, for example, "Exp 10" deviates from linear more than "Exp 8".

### v3

@@ 6.2 Main Results
Table 1 shows that for the low-entropy scenario, watermarked code detection, our detector would achieve the best detection accuracy against other baselines. Specifically, for HumanEval and MBPP respectively, our method could achieve a 3.5% and 9.4% improvement than SWEET in true positive rate while maintaining a low-than 5% false positive rate.
Table 1 shows that for the high-entropy scenarios using Rotowire and C4 datasets, our method is overall very similar to other baselines, sometimes even better than them.
@@ 6.3 Analysis
Impact of Entropy-based Weights. In figure 2, we gave a few visualizations to show the reason why EWD surpassed other baselines in detection performance. Figure 2(a) shows the z-scores of both watermarked and human texts in different datasets, detected by different methods. Our EWD method would result in overall higher z-score for watermarked texts, and slightly lower z-score for human texts. The z-score increase is more significant in low-entropy code generation tasks than in the Rotowire dataset. By enlarging the gap between the z-scores of watermarked and human texts, our detector could better distinguish the watermarked texts. In figure 2(b), we plotted the relationship of a token’s entropy-based weight and probability of that token being sampled from the green list. In watermarked text, the probability is proportional to the token weight. That is to say, the tokens assigned with a large weight are very likely to be sampled from the green list and contribute to a larger z-score. On the contrary, in human-written texts, there doesn’t exist such a relationship, explaining why the z-scores of human text remain low.
Ablation Study. In the experiments, we utilized a linear function to gain the entropy-based weights for each token. We also studied a variety of monotonic increasing functions as shown in figure 3, which are applied in Eq.5 as $f$. Specifically, each function’s gradient is different. For our custom Sigmoid functions, the gradient starts off high and gradually decreases, that is to say, tokens with relatively low entropy will have significant differences in weights, while tokens with relatively high entropy will exhibit smaller differences. Similarly for our custom Exponential functions, the gradient starts off low and gradually increases, meaning that tokens with relatively low entropy will have smaller differences in weight, while tokens with relatively high entropy will exhibit larger differences. We use function strength to represent how far the function deviates from being linear, for example, "Sigmoid 10" deviates from linear more than "Sigmoid 8".
@@
[…]
Table 4: Detection performance of watermarked code using the general prompt instead of original prompts.
Detection w/ General Prompt
Method F1 under 5%FPR Best F1
KGW 0.498 0.741
SWEET 0.580 0.762
EWD 0.600 0.764